Solve the initial value problem below using the method of Laplace transforms. 4ty - 6ty + 6y = 36, y(0) = 6, y(0) = -1 Cl
Answer #1

Q = 36 co solve by laplace transform uty - 6t y + by with yo) = 6, ylo) = -1 som we know that Llyll } s Y(8) - 6 s Lightsa =) y(x) + 2 Y(x) = ogla 2 I order diff eyn which is linean its IF alns e e I 11 2 = is tha soll of 6 6 کرلے + c IF = IF. (required soim. is which S 6 -t (P) h then from o = C => ylo) = 1- o + c So y (O) put world (7) h 7 + 9 8 = = 171R (1) 2 cit